OpenStudy (anonymous):

Compute arc cosecant (1/2)

OpenStudy (jdoe0001):

\(\bf csc^{-1}\left(\frac{1}{2}\right)\implies \cfrac{1}{sin^{-1}\left(\frac{1}{2}\right)}\)

OpenStudy (jdoe0001):

so just use the calculator =)

OpenStudy (jdoe0001):

or the Unit Circle

OpenStudy (anonymous):

So would the answer be pi/6?:)

OpenStudy (jtvatsim):

almost, \[\sin^{-1}(1/2) = \pi/6\]but we want \[\frac{1}{\sin^{-1}(1/2)} = \frac{1}{\pi/6}\]